Suppose that 50 identical batteries are being tested. After 8 hours of continuous use, assume that a given battery is still operating with a probability of 0.70 and has failed with a probability of 0.30.
What is the probability that greater than 30 batteries will last at least 8 hours?

Answer by stanbon(75887)

Suppose that 50 identical batteries are being tested. After 8 hours of continuous use, assume that a given battery is still operating with a
probability of 0.70 and has failed with a probability of 0.30.
What is the probability that greater than 30 batteries will last at least 8 hours?
-----
Binomial Problem with n = 50 and P(op) = 0.0.7
----
P(31<=50) = 1 - P(0<=x<=30) = 1 - binomcdf(50,0.7,30) = 1-0.0848 = 0.9152
----------
Cheers,
Stan H.
----------